OpenStudy (anonymous):

Two forces of magnitude 20N and 40N are acting on a particles such that the angle between them is 135 if both the forces are acting away from the particle calculate their resultant

OpenStudy (dumbcow):

 resultant vector is the sum of the components \[\rightarrow <40\cos135 + 20, 40 \sin 135>\] magnitude: \[\sqrt{(40 \cos135 +20)^2 + (40\sin 135)^2}\]